Perez Corporation runs two convenience stores, one in Connecticut and one in Rhode Island. Operating income for each store in 2017 is as follows:Connecticut Store Rhode Island Store 1,070,000 $ 820,000 Revenues Operating costs Cost of goods sold Lease rent (renewable each year) Labor costs (paid on an hourly basis) Depreciation of equipment Utilities (electricity, heating) Allocated corporate overhead 700,000 87,000 47,000 26,000 40,000 52,000 952,000 118,000 $ 650,000 77,000 42,000 25,000 45,000 37,000 876,000 (56,000) Total operating costs Operating income (loss)1. By closing down the Rhode Island store, Perez can reduce overall corporate overhead costs by $40,000. Calculate Perezs operating income if it closes the Rhode Island store. Is Maria Lopezs statement about the effect of closing the Rhode Island store correct? Explain 2. Calculate Perezs operating income if it keeps the Rhode Island store open and opens another store with revenues and costs identical to the Rhode Island store (including a cost of $25,000 to acquire equipment with a one-year useful life and zero disposal value). Opening this store will increase corporate overhead costs by $7,000. Is Maria Lopezs statement about the effect of adding another store like the Rhode Island store correct? Explain
Answer #1

Answer 1:

Net Income is $ 96,000.

The Rhode Island Store should be closed.

Working note:

Statement of Incremental analysis
Particulars Amount
Revenues $ 1,070,000.00
Less: Expenses
COGS $     700,000.00
Lease rent $       87,000.00
Labor cost $       47,000.00
Depreciation $       51,000.00
Utilities $       40,000.00
Overheads $       49,000.00
Net Income $       96,000.00

Answer 2:

The net income is $ 36,000.

Another store should ot be added.

Statement of Incremental analysis
Particulars Amount
Revenues $    2,710,000.00
Less: Expenses
COGS $    2,000,000.00
Lease rent $       241,000.00
Labor cost $       131,000.00
Depreciation $         76,000.00
Utilities $       130,000.00
Overheads $         96,000.00
Net Income $         36,000.00

Calculation:

Statement of Incremental analysis
Particulars Amount
Revenues =1070000+820000+820000
Less: Expenses
COGS =700000+650000+650000
Lease rent =87000+77000+77000
Labor cost =47000+42000+42000
Depreciation =26000+25000+25000
Utilities =40000+45000+45000
Overheads =52000+37000+7000
Net Income =H75-SUM(H77:H82)

In case of any doubt, please comment.